Occasionally, people can survive falling large distances if the surface they land on is soft enough. During a traverse of Eiger's infamous Nordvand, mountaineer Carlos Ragone's rock anchor gave way and he plummeted 488 feet to land in snow. Amazingly, he suffered only a few bruises and a wrenched shoulder. Assuming that his impact left a hole in the snow 4.5 ft deep, estimate his average acceleration as he slowed to a stop (that is while he was impacting the snow). (Note, for this question the positive direction is defined as down.)
1
Expert's answer
2019-09-02T09:20:03-0400

The velocity just before he touches the snow is given by "v_1"

Now,

using equation of motion for his free fall to snow

"v_1^2-u^2=2a_1s_1"

"u=0"

"a_1=9.8\frac{m}{sec^2}"

"s_1=488ft=148.742m"

Hence,

"v_1=" "54\frac{m}{sec}"

Now,

using equation of motion for decceleration by snow

"v^2-{v_1}^2=2a_2s_2"

"s_2=4.5ft=1.3716m"

"v=0"

Solving it we get,

"a_2=-1063\frac{m}{sec^2}"

his average acceleration as he slowed to a stop is "1063\frac{m}{sec^2}"